OpenStudy (anonymous):

find the values of x at which the curve y = f(x) has a horizontal tangent line. f(x)= (2x + 7)^6 * (x-2)^5

OpenStudy (anonymous):

\[y=(2x+7)^{6}*(x-2)^{5}\] \[y'=5(2x+7)^6*(x-2)^4(1)+6*2(x-2)^5(2x+7)^5\] \[0=5(2x+7)+12(x-2)\] \[0=10x+35+12x-24\] \[22x=-11\] \[x=-1/2\]